OpenStudy (anonymous):

whats the answer of this x+y=10 if x= 7 1/2?

OpenStudy (amistre64):

just plug in the numbers: 7.5 +y = 10; what does y have to be for this to be true?

OpenStudy (amistre64):

7.5 ; if we add .5 we get to 8 if we add 2 more we get to 10 y = 2.5

OpenStudy (amistre64):

or mixed numeralwise 2' 1/2

OpenStudy (anonymous):

if u substitute 7 1/2 to the x is the y become fraction?

OpenStudy (amistre64):

the y will become a fraction yes; it is the value that has to be added to 7' 1/2 to make the equation equal 10

OpenStudy (anonymous):

so whats the graph of x-y=5 and x+y=10